OpenStudy (anonymous):

If sine of x equals square root of 2 over 2, what is cos(x) and tan(x)? Explain your steps in complete sentences.

OpenStudy (anonymous):

sinx=sqrt 2/2 sinx=opp/hyp knowing this, opp=sqrt 2 hyp=2 adj=? (sqrt2)^2+b^2=2^2 , solve for b 2+b^2=4 b^2=2 b= sqrt 2 b is what I called adjacent so know you know adjacent, opposite, and hypotenuse cosx=adj/hyp tanx=opp/adj so cosx=sqrt2/2 tanx=sqrt2/sqrt2=1